OpenStudy (anonymous):

checking for diverging and converging: sum: n=1 -> infinity sin 4n / 4^2 I'm gonna try ratio test a(n+1) / a(n) I think I'm on the right path and get to (sin 4(n+1) / 4) * (1 / sin 4n) I'm not sure how to cancel the sin parts??

OpenStudy (anonymous):

\[\sum_{1}^{\infty} (sin 4n) /(4^{n})\] \[((\sin 4(n+1))/4) * (1/(\sin 4n)\]

OpenStudy (amistre64):

might be better to try an integral test ?

OpenStudy (amistre64):

just a thought, prolly wrong tho

OpenStudy (amistre64):

 aint graphed it yet but im thinking it goes like this

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sin%284n%29%2F4%5En yep

OpenStudy (amistre64):

sin(4(n+1)) and sin(4n) for large values of n are for all practical purposes the same

OpenStudy (amistre64):

and im reading it wrong arent i thats 4^2 underneath or 4^n?

OpenStudy (anonymous):

oh yeah I goofed up the first part its 4^n in the denominator --- is there any manipulation I can do with the sin 4(n+1) / sin 4n or can i jump straight to that is basically a limit of 1 at infinity?

OpenStudy (amistre64):

\[lim\ \frac{sin(4n)}{4^n}*\frac{4^{n-1}}{sin(4n-4) }\] \[lim\ \frac{sin(4n)}{4}*\frac{1}{sin(4n-4)}\] \[\frac 14\ lim\ \frac{sin(4n)}{sin(4n-4)}\] might wanna see if you can split this up with the sin(a+b) identities

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=lim+sin%284n%29%2F4%5En+to+inf or do Lhopital Rule

OpenStudy (amistre64):

sin(4n) doesnt for to inf or 0 tho so Lhop might be out as far as proofing

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=lim+sin%284n%29%2Fsin%284n%2B4%29+to+inf i loathe limits lol

OpenStudy (amistre64):

ratio aint gonna work for you on this

OpenStudy (anonymous):

the limit for sin at infinity would be one in any case, ya? which makes sense with the wolfram answer 1/really big number = 0

OpenStudy (amistre64):

i dont know what the wolfs basing it on; but i know that the sin will occilate back and forth inbetween 1/4^n which goes to zero

OpenStudy (amistre64):

4n^(-1) is the amplitude of sin; and as n goes larg; 4^n goes small and approaches zero

OpenStudy (amistre64):

the squeeze thrm would say that sin gets squaeezed to zero along with it

OpenStudy (amistre64):